# Spatial-Spectral Morphological Mamba for Hyperspectral Image Classification

### v1

@@ VI Ablation Study and Discussion
VI-A Comparison of Spatial-Spectral Mamba With and Without Morphology
@@ VI-B Feature Spaces
t-Distributed Stochastic Neighbor Embedding (t-SNE) is an effective method for visualizing high-dimensional HSI data [58]. t-SNE operates in an unsupervised, non-linear manner to preserve pairwise similarities between data points in a lower-dimensional space, focusing on maintaining small pairwise distances. It achieves this by first determining similarity measures between pairs of instances in both higher and lower-dimensional spaces [59, 60], then optimizing these measures through gradient descent. t-SNE models the likelihood of one point being a neighbor of another in both high and low dimensions. It computes pairwise similarities in the high-dimensional space using a Gaussian kernel, mapping these points onto a lower-dimensional space while preserving their similarities. This is achieved by minimizing the divergence between the probability distributions of high and low-dimensional spaces. The process results in clusters of similar data points in the lower-dimensional space, aiding in the visualization and understanding of high-dimensional data. Figures 13(a), 13(b), 13(c), 13(d), and 13(e) show the feature representations learned by the MorpMamba model for the PU, PC, SA, UH, and LK datasets, respectively, highlighting its effectiveness in feature learning for classification.
@@ VII Conclusions and Future Research Directions
MorpMamba introduces a novel approach to HSIC by integrating spatial-spectral morphological operations with the Mamba architecture. The model has demonstrated SOTA performance across multiple publicly available HS datasets, achieving superior results in terms of overall accuracy, average accuracy, and Kappa ($\kappa$) coefficient. Additionally, MorpMamba shows remarkable computational efficiency, utilizing significantly fewer parameters while maintaining or surpassing the classification accuracy of other SOTA models. The integration of morphological operations (erosion and dilation) in the tokenization process enables the model to capture both fine details and global structures in spatial and spectral dimensions. Moreover, the token enhancement module dynamically adjusts the importance of spatial and spectral features, leading to more robust and context-aware feature representation. The multi-head self-attention mechanism effectively captures complex relationships and dependencies within the HS data.
@@
Further optimization is needed to robustly handle high-dimensional HS data, focusing on more efficient data processing techniques tailored for HS imagery. To enhance the model’s ability to generalize across diverse datasets, strategies such as transfer learning, domain adaptation, or meta-learning should be explored. Integrating advanced feature extraction techniques, like graph neural networks or novel attention-based mechanisms, could improve the model’s ability to capture complex spatial-spectral relationships. Comprehensive evaluations on larger and more diverse HS datasets are necessary to validate the model’s performance and identify areas for improvement, including testing on datasets from various geographical regions, sensor types, and challenging environmental conditions. Developing hardware-aware optimizations and deployment strategies for MorpMamba could enable its use on resource-constrained devices, allowing for on-board processing in satellite or drone-based systems. Enhancing the interpretability of the model’s decisions by visualizing the importance of different spectral bands or spatial regions could provide valuable insights for domain experts and increase trust in the model’s outputs. Exploring the potential of MorpMamba for multi-task learning scenarios, such as simultaneous classification and segmentation of HS images, could maximize the utility of the extracted features. Investigating the application of MorpMamba to multi-temporal HS data could leverage its ability to model the temporal evolution of features for change detection or time-series analysis of land cover changes. Additionally, integrating MorpMamba with other remote sensing data sources, such as LiDAR or SAR, could create more comprehensive and accurate Earth observation models. By addressing these future directions, the research community can build upon the foundations laid by MorpMamba to further advance HSI analysis and its applications in remote sensing and Earth observation.

### v2

@@ VI Ablation Study and Discussion
t-Distributed Stochastic Neighbor Embedding (t-SNE) is an effective method for visualizing high-dimensional HSI data. t-SNE operates in an unsupervised, non-linear manner to preserve pairwise similarities between data points in a lower-dimensional space, focusing on maintaining small pairwise distances. It achieves this by first determining similarity measures between pairs of instances in both higher and lower-dimensional spaces [56], then optimizing these measures through gradient descent. t-SNE models the likelihood of one point being a neighbor of another in both high and low dimensions. It computes pairwise similarities in the high-dimensional space using a Gaussian kernel, mapping these points onto a lower-dimensional space while preserving their similarities. This is achieved by minimizing the divergence between the probability distributions of high and low-dimensional spaces. The process results in clusters of similar data points in the lower-dimensional space, aiding in the visualization and understanding of high-dimensional data. Figures 12(a), 12(b), 12(c), 12(d), and 12(e) show the feature representations learned by the MorpMamba model for the PU, PC, SA, UH, and LK datasets, respectively, highlighting its effectiveness in feature learning for classification.
@@ VII Conclusions and Future Research Directions
MorpMamba introduces a novel approach to HSI classification by integrating spatial-spectral morphological operations with the Mamba architecture. The model has demonstrated SOTA performance across multiple publicly available HSI datasets, achieving superior results in terms of overall accuracy, average accuracy, and Kappa ($\kappa$) coefficient. Additionally, MorpMamba shows remarkable computational efficiency, utilizing significantly fewer parameters while maintaining or surpassing the classification accuracy of other SOTA models. The integration of morphological operations (erosion and dilation) in the tokenization process enables the model to capture both fine details and global structures in spatial and spectral dimensions. Moreover, the token enhancement module dynamically adjusts the importance of spatial and spectral features, leading to more robust and context-aware feature representation. The multi-head self-attention mechanism effectively captures complex relationships and dependencies within the HSI data.
@@
Further optimization is needed to robustly handle high-dimensional HSI data, focusing on more efficient data processing techniques tailored for HSI imagery. To enhance the model’s ability to generalize across diverse datasets, strategies such as transfer learning, domain adaptation, or meta-learning should be explored. Integrating advanced feature extraction techniques, like graph neural networks or novel attention-based mechanisms, could improve the model’s ability to capture complex spatial-spectral relationships. Comprehensive evaluations on larger and more diverse HSI datasets are necessary to validate the model’s performance and identify areas for improvement, including testing on datasets from various geographical regions, sensor types, and challenging environmental conditions. Developing hardware-aware optimizations and deployment strategies for MorpMamba could enable its use on resource-constrained devices, allowing for on-board processing in satellite or drone-based systems. Enhancing the interpretability of the model’s decisions by visualizing the importance of different spectral bands or spatial regions could provide valuable insights for domain experts and increase trust in the model’s outputs. Exploring the potential of MorpMamba for multi-task learning scenarios, such as simultaneous classification and segmentation of HSI images, could maximize the utility of the extracted features. Investigating the application of MorpMamba to multi-temporal HSI data could leverage its ability to model the temporal evolution of features for change detection or time-series analysis of land cover changes. Additionally, integrating MorpMamba with other remote sensing data sources, such as LiDAR or SAR, could create more comprehensive and accurate Earth observation models. By addressing these future directions, the research community can build upon the foundations laid by MorpMamba to further advance HSI analysis and its applications in remote sensing and Earth observation.
Acknowledgment
The authors extend their appreciation to the Researchers Supporting Project number (RSPD2024R848), King Saud University, Riyadh, Saudi Arabia.